# Repairing Reward Functions with Human Feedback to Mitigate Reward Hacking

### v1

@@ 4 Methodology
The first term, $\mathcal{L}_{\text{pref}}$, is a standard preference loss that encourages the modified reward function $\hat{r}_{\text{proxy}}+g$ to satisfy the preferences in $\mathcal{D}_{t}$. The second term $\mathcal{L}^{+}$ regularizes the correction term $g$ towards zero on trajectory pairs where the proxy reward function agrees with the human preference. This assumes that when the proxy reward function correctly ranks a pair of trajectories, its assigned reward for the transitions in those trajectories is consistent with the ground-truth reward function; $\mathcal{L}^{+}$ prevents unnecessary adjustments that could degrade an otherwise correct reward signal. Finally, the third term $\mathcal{L}^{-}$ focuses on trajectory pairs that were misclassified by the modified reward function. Adjusting the correction term $g$ to correctly classify such trajectories is generally underspecified– one could (a) increase the reward for the preferred trajectory, or (b) decrease the reward for the not preferred trajectory. Consistent with our assumptions on the proxy reward function, $\mathcal{L}^{-}$ prioritizes option (b), regularizing the correction term $g$ to zero on transitions in the preferred trajectories, which will prioritize a negative correction for undesirable behaviors.
To ensure our approach still learns the ground-truth reward function when our assumptions about the specified proxy reward function fail to hold, we decay $\lambda_{1}$ and $\lambda_{2}$ over iterations (see Appendix E.6).
@@ 4.1 Reward Hacking Analysis
We begin with the following assumptions, which differ from those in Section 4:
@@ 5.2 Baselines
We compare PBRR against state-of-the-art and natural ablation baselines. More details are in Appendix D. Appendix K.3 presents additional experiments where we update the reference policy across iterations, when relevant.
@@ 5.3 Main results: repairing $\hat{r}_{\text{proxy}}$ with preferences
Outperforming the reference policy $\pi_{\text{ref}}$    Our theoretical analysis of PBRR with $C_{1}=0$ (Appendix J) only guarantees that it asymptotically performs no worse than the reference policy. See Appendix H for illustrative examples. Nevertheless, PBRR empirically induced policies that significantly outperform the reference policy in all environments.
@@
We also sought to isolate the contributions of our preference-learning objective (Eq. 4) and exploration strategy (Section 4). Figure 3 shows that repairing the proxy reward function with PBRR but using the standard loss in Eq. 1 instead of Eq.  4 yields substantially less stable performance and a lower mean-return under the ground-truth reward function across all environments. Without $\mathcal{L}^{+}$ and $\mathcal{L}^{-}$ from Eq. 4, the updated proxy reward function incorrectly assigns a higher reward to the suboptimal actions taken by the reference policy. Appendix G.3 provides a qualitative analysis in the AI safety gridworld, and Appendix K.4 reports further ablations of each regularization term. Conversely, using PBRR’s objective (Eq. 4) within the RRM or RRM+State-Constraint baselines fails to match PBRR’s data efficiency in all environments except Traffic Control. This shows that both PBRR’s preference-learning objective and exploration strategy are important to efficiently repair the proxy reward function with preferences.
@@ 6 Conclusion
We introduce Preference-Based Reward Repair (PBRR), a framework to repair a human-specified proxy reward function by learning a transition-level correction from preferences. Across a diverse set of benchmarks, spanning autonomous vehicle traffic control, pandemic lockdown regulation design, and insulin regulation for diabetes patients, PBRR achieves higher performance with greater stability and fewer preferences than methods that either learn a reward function from scratch or modify the proxy reward function using alternative strategies. Our ablations show that both components of PBRR are necessary: the exploration strategy that leverages a supplied reference policy to identify transitions over which the proxy reward function is incorrect, and the preference-learning objective that encourages only correcting the proxy reward function on transitions where it incorrectly assigns high reward. Our theoretical analysis further shows that a variant of PBRR attains a sub-linear regret bound comparable with prior work. Additional techniques for improving RLHF data efficiency—such as data augmentation (Park et al., 2022) and intrinsic exploration rewards (Liang et al., 2022)—remain untested in our domains and are orthogonal to our contributions. Future work could explore combining these methods with PBRR to see if further gains in data efficiency are possible. Overall, our results suggest that repairing a human-specified proxy reward function with PBRR, rather than learning a reward function from scratch, is a data-efficient path to alignment in complex, sequential decision making tasks.
@@
Contribution Statement Stephane Hatgis-Kessell led the ideation, methodology development, and experimental implementation. Logan Mondal Bhamidipaty led the theoretical analysis of reward hacking in Section 4.1 and contributed to methodology development. Emma Brunskill led the theoretical analysis of the regret bounds in Section 4 and provided extensive guidance on methodology design and experimentation. All three authors wrote the paper.
@@ Appendix B Environment Details
Appendix C Optimizing for $\hat{r}$ to avoid reward hacking
@@ Appendix I Regret Bound Proofs
Here we provide a theoretical analysis motivating PBRR, relying on different assumptions than the regret analysis in Section 4. We show that PBRR’s repaired reward function is guaranteed to induce an optimal policy that matches or exceeds the performance of the reference policy (Thm. 4.3) and resolve two specific instantiations of reward hacking (Cors. 4.4, 4.5) in the infinite data setting as the number of iterations goes to infinity.

### v2

# Repairing Reward Functions with Feedback to Mitigate Reward Hacking

@@ 4 Methodology
The first term, $\mathcal{L}_{\text{pref}}$, is a standard preference loss that encourages the modified reward function $\hat{r}_{\text{proxy}}+g$ to satisfy the preferences in $\mathcal{D}_{t}$. The second term $\mathcal{L}^{+}$ regularizes the correction term $g$ towards zero on trajectory pairs where the proxy reward function agrees with the human preference. This assumes that when the proxy reward function correctly ranks a pair of trajectories, its assigned reward for the transitions in those trajectories is consistent with the ground-truth reward function; $\mathcal{L}^{+}$ prevents unnecessary adjustments that could degrade an otherwise correct reward signal. Finally, the third term $\mathcal{L}^{-}$ focuses on trajectory pairs that were misclassified by the modified reward function. Adjusting the correction term $g$ to correctly classify such trajectories is generally underspecified– one could (a) increase the reward for the preferred trajectory, or (b) decrease the reward for the not preferred trajectory. $\mathcal{L}^{-}$ prioritizes option (b), regularizing the correction term $g$ to zero on transitions in the preferred trajectories, which will prioritize a negative correction for undesirable behaviors.
To ensure our approach still learns the ground-truth reward function when our assumptions about the specified proxy reward function fail to hold, we decay $\lambda_{1}$ and $\lambda_{2}$ over iterations (see Appendix E.6). Although Eq. 4 leverages the assumption that the proxy reward function is optimistic, our algorithm does not require this assumption, nor does our theoretical analysis in Section 4.1 or Section 4.2.
@@ 4.2 Reward Hacking Analysis
We begin with the following assumptions, which differ from those in Section 4.1:
@@ 5.2 Baselines
We compare PBRR against state-of-the-art and natural ablation baselines. More details are in Appendix D.
@@ 5.3 Main results: repairing $\hat{r}_{\text{proxy}}$ with preferences
Outperforming the reference policy $\pi_{\text{ref}}$    PBRR induced policies that outperform the reference policy in all environments. The reference policy used by PBRR need not be performant; it must only provide a useful comparison to the proxy reward function’s induced policy. In Appendix K.8, we empirically show that PBRR continues to match or outperform all baselines when using a randomly initialized reference policy.
Optimism Assumption    PBRR does not require the input reward function to be optimistic—note the decay parameters in Algorithm 1—though it may do particularly well in such a setting. We explore the performance PBRR when the input reward function is not optimistic in two domains, and find that PBRR continues to be the most performant method. In the Glucose Monitoring environment, the optimism assumption does not hold; policies that maximize patient health outcomes, i.e., are optimal under the ground-truth reward function, are penalized under the proxy reward function due to their financial cost. PBRR still outperforms all baselines. Appendix K.7 reports similar findings in the AI Safety Gridworld when repairing a proxy reward function that is not optimistic.
@@
We also sought to isolate the contributions of our preference-learning objective (Eq. 4) and exploration strategy (outlined in Algorithm 1). Figure 3 shows that repairing the proxy reward function with PBRR but using the standard loss in Eq. 1 instead of Eq.  4 yields substantially less stable performance and a lower mean-return under the ground-truth reward function across all environments. Without $\mathcal{L}^{+}$ and $\mathcal{L}^{-}$ from Eq. 4, the updated proxy reward function incorrectly assigns a higher reward to the suboptimal actions taken by the reference policy. Appendix G.3 provides a qualitative analysis in the AI safety gridworld, and Appendix K.4 reports further ablations of each regularization term. Conversely, using PBRR’s objective (Eq. 4) within the RRM or RRM+State-Constraint baselines fails to match PBRR’s data efficiency in all environments except Traffic Control. This shows that both PBRR’s preference-learning objective and exploration strategy are important to efficiently repair the proxy reward function with preferences.
@@ 6 Conclusion
We introduce Preference-Based Reward Repair (PBRR), a framework to repair a human-specified proxy reward function by learning a transition-level correction from preferences. Across a diverse set of benchmarks, spanning autonomous vehicle traffic control, pandemic lockdown regulation design, and insulin regulation for diabetes patients, PBRR achieves higher performance with greater stability and fewer preferences than methods that either learn a reward function from scratch or modify the proxy reward function using alternative strategies. Our ablations show that both components of PBRR are necessary: the exploration strategy that leverages a supplied reference policy to identify transitions over which the proxy reward function is incorrect, and the preference-learning objective that encourages only correcting the proxy reward function on transitions where it incorrectly assigns high reward. Our theoretical analysis further shows that a variant of PBRR attains a sub-linear regret bound comparable with prior work. Overall, our results suggest that repairing a human-specified proxy reward function with PBRR, rather than learning a reward function from scratch, is a data-efficient path to alignment in complex, sequential decision making tasks.
@@
Contribution Statement Stephane Hatgis-Kessell led the ideation, methodology development, and experimental implementation. Logan Mondal Bhamidipaty led the theoretical analysis of reward hacking in Section 4.2 and contributed to methodology development. Emma Brunskill led the theoretical analysis of the regret bounds in Section 4.1 and provided extensive guidance on methodology design and experimentation. All three authors wrote the paper.
@@ Appendix B Environment Details
Appendix C Optimizing for $\hat{r}$ while avoiding reward hacking
@@ Appendix I Regret Bound Proofs
Here we provide a theoretical analysis motivating PBRR, relying on different assumptions than the regret analysis in Section 4.1. We show that PBRR’s repaired reward function is guaranteed to induce an optimal policy that matches or exceeds the performance of the reference policy (Thm. 4.3) and resolve two specific instantiations of reward hacking (Cors. 4.4, 4.5) in the infinite data setting as the number of iterations goes to infinity.
@@
K.8 PBRR with a randomly initialized reference policy
PBRR requires a reference policy for preference elicitation. For all prior experiments, unless otherwise stated, we use a realistic reference policy that a human could provide such as one trained with a handful of human demonstrations or from imperfect hand-written rules. Here, we investigate PBRR’s performance when using a randomly initialize reference policy instead. Figure 15 shows that performance is largely unchanged; the quality of the reference policy used by PBRR is not judged by its performance, but rather its coverage of the state-action space relative to the policy induced by the proxy reward function. These results imply that for many tasks, a sufficient reference policy should always be available by simply using a randomly initialized policy.
Figure 15: Mean return under the ground-truth reward function achieved by PBRR with the default reference policy used for all other experiments (PBRR ($\pi_{\text{ref}}-\text{default}$)), compared to (i) PBRR using a randomly initialized reference policy (PBRR ($\pi_{\text{ref}}-\text{random-init}$)), and (ii) the Online-RLHF method. The plot for the Pandemic Mitigation environment is truncated to make it easier to differentiate performance between various methods; the full plot is shown in Figure 16.
Figure 16: Mean return under the ground-truth reward function achieved by PBRR with the default reference policy used for all other experiments (PBRR ($\pi_{\text{ref}}-\text{default}$)), compared to (i) PBRR using a randomly initialized reference policy (PBRR ($\pi_{\text{ref}}-\text{random-init}$)), and (ii) the Online-RLHF method. For easier visual comparison, we truncate the plot for the Pandemic Mitigation environment in Figure 15.